# THEME : Enhancing Thematic Investing with Semantic Stock Representations and Temporal Dynamics

### v1

@@ Abstract.
Thematic investing aims to construct portfolios aligned with structural trends, yet selecting relevant stocks remains challenging due to overlapping sector boundaries and evolving market dynamics. To address this challenge, we construct the Thematic Representation Set (TRS), an extended dataset that begins with real-world thematic ETFs and expands upon them by incorporating industry classifications and financial news to overcome their coverage limitations. The final dataset contains both the explicit mapping of themes to their constituent stocks and the rich textual profiles for each. Building on this dataset, we introduce THEME, a hierarchical contrastive learning framework. By representing the textual profiles of themes and stocks as embeddings, THEME first leverages their hierarchical relationship to achieve semantic alignment. Subsequently, it refines these semantic embeddings through a temporal refinement stage that incorporates individual stock returns. The final stock representations are designed for effective retrieval of thematically aligned assets with strong return potential. Empirical results show that THEME outperforms strong baselines across multiple retrieval metrics and significantly improves performance in portfolio construction. By jointly modeling thematic relationships from text and market dynamics from returns, THEME provides a scalable and adaptive solution for navigating complex investment themes.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/XXXXXX.XXXXXXccs: Applied computing Economicsccs: Information systems Information retrieval
@@ 1. Introduction
Figure 1. Comparison of t-SNE visualizations of stock description embeddings. (Upper Left) A baseline embedding model fails to cluster stocks by thematic meaning. (Lower Right) Our tuned embeddings organize stocks by economic theme, enabling more interpretable and structured retrieval.
@@
  • A domain-tuned embedding space for finance: We show that general-purpose embeddings fail to organize stocks by thematic relevance. In contrast, our fine-tuned embeddings yield clearer semantic structure, improving both interpretability and retrieval quality.
@@ 3.2. Stage 2: Temporal Refinement
The second stage refines the semantic embeddings by incorporating dynamic, short-term return signals. To achieve this, we introduce a lightweight 2-layer adapter, denoted as $\mathcal{A}_{\phi}$, with trainable parameters $\phi$. This adapter is designed to fuse semantic context with dynamic temporal patterns. It takes two distinct inputs for a stock $s_{j}$: its semantic embedding ${h}_{j}$ from Stage 1, and its past $L$ trading days of daily returns, ${r}_{j}$, where $L$ is the lookback period. The adapter jointly processes these inputs to produce a single, temporally-refined stock embedding ${h}^{\prime}_{j}$:
@@
The adapter’s parameters $\phi$ are optimized via a triplet loss function designed to rank stocks within a theme based on future returns. For a given theme $t_{i}$, we first select a positive sample $s_{p}$ and a negative sample $s_{n}$ from its constituent stocks. The stock with the higher forward return over a horizon of $H$ days is designated as the positive sample, while the other serves as the negative. These samples, along with the theme’s semantic embedding $z_{i}$ as an anchor, form a training triplet $(z_{i},h^{\prime}_{p},h^{\prime}_{n})$, where $h^{\prime}_{p}$ and $h^{\prime}_{n}$ are the temporally-refined embeddings of the positive and negative stocks, respectively. The training objective is to minimize the following loss, thereby ensuring that the anchor is closer in the embedding space to the positive sample than to the negative sample by at least a margin $m$:
@@ 3.4. System Implementation and Integration
[…]
@@ 4.2. Experiment 1: Retrieval Performance
We compare THEME against multiple baselines including embedding models (e.g., Fin-E5, SFR-Embedding-Mistral), open LLMs (GPT-4.1, Gemini-2.5), and our own Linq-Embed-Mistral. Each model ranks a shared stock universe given a theme description, and we evaluate retrieval quality at $k{=}\{3,5,10\}$ using HR and P.
As shown in Table 1, THEME significantly improves retrieval quality across all model backbones. For example, when applied to Linq-Embed-Mistral, THEME improves HR@3 from 0.5155 to 0.8196 and P@3 from 0.3522 to 0.6289. Similar gains are observed across other 7B-scale models such as SFR and e5-mistral. Even smaller base models (e.g., bge-small-en-v1.5, multilingual-e5) show substantial performance boosts when integrated with THEME, often exceeding the retrieval quality of much larger models in their vanilla form.
Notably, THEME-enhanced versions consistently outperform strong LLM baselines such as GPT-4.1 and Gemini-2.5, which are not directly compatible with contrastive fine-tuning. This highlights the benefit of incorporating theme-aligned supervision at scale. Overall, these results confirm that THEME is effective at improving thematic relevance in stock retrieval tasks, regardless of the underlying model size.
@@ 4.3. Experiment 2: Portfolio Construction
Building on the results from Experiment 4.2, which demonstrated that THEME effectively retrieves thematically relevant stocks, we now assess whether these stocks also yield superior investment performance. To this end, we evaluate the performance of portfolios constructed using the top-$K$ stocks selected by THEME over a rolling test period from April 23, 2024 to April 29, 2025.
For each window, we construct equal-weighted portfolios using the top-$K$ stocks ($K\in\{3,5,10\}$) ranked by similarity to the theme anchor. The average daily return of these $K$ stocks is then recorded over the subsequent 14 trading days. Chaining the returns across all windows produces a continuous daily return series for the entire test period.
@@
As summarized in Table 2, THEME consistently improves portfolio quality over vanilla approaches. For example, with the gte-Qwen2-7B-instruct model, SR@3 increases from 0.5014 to 0.7592, CR@3 from 0.0917 to 0.1645, and MDD improves from –0.2427 to –0.2378. These results also outperform the baseline performance of investing in the real thematic ETF universe, which achieves a Sharpe ratio of 0.4845, CR of 0.0672, and MDD of –0.2368. Similar advantages are observed across other $K$ values and model variants.
Even with strong baselines such as GritLM-7B, THEME maintains or slightly improves profitability metrics while remaining robust across architectures. These findings indicate that combining semantic and temporal signals enables the retrieval of stocks that are not only thematically aligned but also effective in constructing more profitable and resilient portfolios in dynamic markets.
@@ 5. Discussion
Third, our ablation study in Experiment 4.4 emphasizes two design decisions that directly impact real-world performance: the choice of semantic anchors and the composition of the training dataset. We find that training with abstract theme descriptions as anchors yields better generalization than stock-to-stock alignment. In parallel, training on the expanded Thematic Representation Set (TRS) leads to consistent retrieval gains over using real ETF data alone. This is especially important for practitioners seeking to cover a wide array of investment themes, including those not yet captured in the current ETF landscape. TRS also enables multi-theme associations, ensuring that stocks with diverse business models are not artificially constrained to a single category.
@@
Looking forward, we are incorporating real-time data sources, such as ESG events, earnings calls, and patent filings, to support additional use cases like intraday thematic scoring, event-driven rebalancing, or early detection of theme dislocation. Rather than replacing existing approaches, THEME serves as a thematic intelligence layer that augments decision-making across the spectrum of passive and active strategies.
Overall, THEME bridges the gap between qualitative narratives and quantitative execution. By translating loosely defined thematic concepts into structured, time-sensitive representations, it enables more adaptive and scalable participation in evolving markets.
@@ 6. Conclusion
This work makes three key contributions to thematic investing. First, we propose THEME, a hierarchical contrastive learning framework that integrates semantic and temporal supervision to produce theme-aware stock embeddings. Second, we construct the TRS, a scalable dataset that addresses the coverage limitations of real-world thematic ETFs by incorporating industry classifications and financial news. Third, we demonstrate through extensive experiments that these design choices lead to consistent improvements in retrieval accuracy and portfolio performance across a range of language model backbones, while supporting real-world deployment through modular APIs and customizable interfaces.
@@
Beyond performance, THEME offers practical utility across investment workflows. Its outputs support thematic universe construction, dynamic portfolio rebalancing, and personalized investing. The system’s modular, API-based architecture enables integration into discretionary platforms, systematic engines, or client-facing applications.
@@
Overall, THEME shows how semantic understanding and temporal modeling can be jointly applied to build more adaptive, scalable, and insightful systems for navigating complex investment themes in modern financial markets.
@@ Acknowledgments
This work was supported by National Research Foundation of Korea (NRF) grant (No. NRF-2022R1I1A4069163) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant (No. RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)) funded by the Korea government(MSIT). We would like to thank LG AI Research for providing computational and research support.

### v2

@@ Abstract.
Thematic investing, which aims to construct portfolios aligned with structural trends, remains a challenging endeavor due to overlapping sector boundaries and evolving market dynamics. A promising direction is to build semantic representations of investment themes from textual data. However, despite their power, general-purpose LLM embedding models are not well-suited to capture the nuanced characteristics of financial assets, since the semantic representation of investment assets may differ fundamentally from that of general financial text. To address this, we introduce THEME, a framework that fine-tunes embeddings using hierarchical contrastive learning. THEME aligns themes and their constituent stocks using their hierarchical relationship, and subsequently refines these embeddings by incorporating stock returns. This process yields representations effective for retrieving thematically aligned assets with strong return potential. Empirical results demonstrate that THEME excels in two key areas. For thematic asset retrieval, it significantly outperforms leading large language models. Furthermore, its constructed portfolios demonstrate compelling performance. By jointly modeling thematic relationships from text and market dynamics from returns, THEME generates stock embeddings specifically tailored for a wide range of practical investment applications.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761517ccs: Applied computing Economicsccs: Information systems Information retrieval
@@ 1. Introduction
Figure 1. A t-SNE comparison of stock embeddings before and after applying THEME. (Upper Left) The baseline model fails to group stocks by investment theme. (Lower Right) In contrast, our tuned embeddings form distinct, meaningful clusters corresponding to investment themes like FinTech and Future Vehicles, significantly improving interpretability.
@@
  • A domain-tuned embedding space for finance: We show that general-purpose embeddings fail to organize stocks by thematic relevance. In contrast, our fine-tuned embeddings yield clearer semantic structure, improving retrieval and interpretability.
@@ 3.2. Stage 2: Temporal Refinement
The second stage refines the semantic embeddings by incorporating dynamic, short-term return signals. To achieve this, we introduce a lightweight 2-layer adapter, denoted as $\mathcal{A}_{\phi}$, with trainable parameters $\phi$. This adapter is designed to fuse semantic context with dynamic temporal patterns. It takes two distinct inputs for a stock $s_{j}$: its semantic embedding ${h}_{j}$ from Stage 1, and its past $L$ trading days of daily returns, ${r}_{j}$, where $L$ is the lookback period. The adapter jointly processes these inputs to produce a single, fusion embedding ${h}^{\prime}_{j}$:
@@
The adapter’s parameters $\phi$ are optimized via a triplet loss function designed to rank stocks within a theme based on future returns. For a given theme $t_{i}$, we first select a positive sample $s_{p}$ and a negative sample $s_{n}$ from its constituent stocks. The stock with the higher forward return over a horizon of $H$ days is designated as the positive sample, while the other serves as the negative. These samples, along with the theme’s semantic embedding $z_{i}$ as an anchor, form a training triplet $(z_{i},h^{\prime}_{p},h^{\prime}_{n})$, where $h^{\prime}_{p}$ and $h^{\prime}_{n}$ are the fusion embeddings of the positive and negative stocks, respectively. The training objective is to minimize the following loss, thereby ensuring that the anchor is closer in the embedding space to the positive sample than to the negative sample by at least a margin $m$:
@@ 3.4. System Implementation and Integration
voyage-2-finance (VoyageAI, 2025) Unknown 0.4278 - 0.2285 - 0.4742 - 0.2134 - 0.5567 - 0.1948 -
Fin-E5 (Tang and Yang, 2025) Unknown 0.4948 - 0.3247 - 0.5515 - 0.3113 - 0.6289 - 0.2835 -
GPT-4.1 (OpenAI, 2023) Unknown 0.7113 - 0.5189 - 0.7731 - 0.4505 - 0.8092 - 0.3690 -
Gemini-2.5 (Gemini Team, 2025) Unknown 0.6494 - 0.4020 - 0.6649 - 0.3463 - 0.7268 - 0.3190 -
@@ 4.2. Experiment 1: Retrieval Performance
In this experiment, we evaluate the retrieval performance of THEME against various base models. The evaluation includes several types of models: domain-specific embedding models (gray rows), state-of-the-art LLMs (yellow rows), and general-purpose embedding models. Each model ranks a universe of stocks based on a given theme description, and we measure the retrieval quality using Hit Rate (HR) and Precision (P) at $k\in\{3,5,10\}$.
The results in Table 1 clearly show that THEME dramatically improves retrieval performance. For example, applying THEME to our Linq-Embed-Mistral model boosted its HR@3 from 0.5155 to 0.8196 and P@3 from 0.3522 to 0.6289. Similar gains were observed across other 7B-scale models. Notably, this enhancement was not limited to large models. Even the small bge-small-en-v1.5 model, when combined with THEME, surpassed larger vanilla 7B models and the domain-specific embedding models.
Furthermore, models enhanced with THEME recorded superior performance compared to powerful SOTA LLMs such as GPT-4.1 and Gemini-2.5. This highlights that while general-purpose LLMs possess a vast amount of knowledge, a more specialized strategy of integrating theme-aligned supervision is ultimately more effective for this specific task. Ultimately, the top performance of models like Linq-Embed-Mistral and GritLM-7B confirms that THEME is an effective and model-agnostic methodology for significantly enhancing thematic stock retrieval, regardless of the underlying model’s size or type.
@@ 4.3. Experiment 2: Portfolio Construction
Having confirmed that THEME retrieves relevant stocks (Experiment 4.2), we now assess their investment performance. To this end, we evaluate the performance of portfolios constructed using the top-$K$ stocks selected by THEME over a rolling test period from April 23, 2024 to April 29, 2025.
For each window, we construct equal-weighted portfolios using the top-$K$ stocks ranked by similarity to the theme anchor. The average daily return of these $K$ stocks is then recorded over the subsequent 14 trading days. Chaining the returns across all windows forms a continuous daily return series over the entire test period.
@@
As summarized in Table 2, THEME consistently improves portfolio quality over vanilla approaches. The performance gains are clear: for the gte-Qwen2-7B-instruct, SR@3 increases from 0.5014 to 0.7592, CR@3 from 0.0917 to 0.1645, and MDD improves from –0.2427 to –0.2378. These results also outperform the baseline performance of investing in the real thematic ETFs, which achieves a Sharpe ratio of 0.4845, CR of 0.0672, and MDD of –0.2368. Similar advantages are observed across other $K$ values and model variants.
Notably, the impact of THEME varies depending on the baseline’s initial strength. On a powerful baseline like GritLM-7B, which performs exceptionally well in its vanilla state, our method offers marginal gains in specific top-3 scenarios while not consistently surpassing it across all metrics. This suggests that our temporal refinement provides the most significant lift to models that are not yet fully optimized for the task, such as gte-Qwen2-7B-instruct. Ultimately, these findings confirm that fusing semantic and temporal signals is a highly effective strategy for creating profitable and resilient portfolios.
@@ 5. Discussion
Third, our ablation study in Experiment 4.4 emphasizes two design decisions that directly impact real-world performance: the choice of semantic anchors and the composition of the training dataset. We find that training with abstract theme descriptions as anchors yields better generalization than stock-to-stock alignment. In parallel, training on the expanded TRS leads to consistent retrieval gains over using real ETF data alone. This is especially important for practitioners seeking to cover a wide array of investment themes, including those not yet captured in the current ETF landscape. TRS also enables multi-theme associations, ensuring that stocks with diverse business models are not artificially constrained to a single category.
@@
Looking forward, we are incorporating real-time data sources, such as ESG events, earnings calls, and patent filings, to support additional use cases like intraday thematic scoring or event-driven rebalancing. Rather than replacing existing approaches, THEME is designed to serve as a thematic intelligence layer that augments decision-making for both passive and active strategies.
@@ 6. Conclusion
This work makes three key contributions to thematic investing. First, we propose THEME, a hierarchical contrastive learning framework that integrates semantic and temporal supervision to produce theme-aware stock embeddings. Second, we construct the TRS, a scalable dataset that addresses the coverage limitations of real-world thematic ETFs by incorporating industry classifications and financial news. Third, we demonstrate through extensive experiments that these design choices lead to consistent improvements in retrieval accuracy and portfolio performance, yielding a practical framework where embeddings can be pre-computed and stored for on-demand use.
@@
Beyond performance, THEME offers practical utility across investment workflows. Its outputs support thematic universe construction, dynamic portfolio rebalancing, and personalized investing. The resulting theme-aware embeddings are structured for efficient retrieval, enabling their integration into discretionary platforms, systematic engines, or client-facing applications.
@@
In essence, THEME shows how semantic understanding and temporal modeling can be jointly applied to build more adaptive, scalable, and insightful systems for navigating complex investment themes in modern financial markets.
@@ Acknowledgments
This work was supported by National Research Foundation of Korea (NRF) grant (No. NRF-2022R1I1A4069163) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant (No. RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)) funded by the Korea government(MSIT).